If a and b are roots of 7x^2 − 2002x +182 = 0 , find the value of 1/a + 1/b
Found 2 solutions by Alan3354, JimboP1977:
Answer by Alan3354(69443)

If a and b are roots of 7x^2 − 2002x +182 = 0 , find the value of 1/a + 1/b
------------------
7x^2 − 2002x +182 = 0
x^2 - 286x + 26 = 0
The roots are:
(-b + sqrt(b^2 - 4ac))/2a and
(-b - sqrt(b^2 - 4ac))/2a
----------------------------
The inverses give:
2a/(-b + sqrt(b^2 - 4ac)) and
2a/(-b - sqrt(b^2 - 4ac))
The sum is:
2a*2*(-b)/(b^2 - (b^2 - 4ac)
= 4ab/(4ac)
= b/c The general solution
-----------------
For this eqn it's -11
= -11

Answer by JimboP1977(311)

I made 1/a+1/b= 11 (positive)